Hi. Please explain me why the answer right and wrong. Make it clear please, thanks!
Question 9 0 out of 1 points On anuary 1 Year 1 Stratton Company borrowed $140,000 on a 10-year 6% installment note payable. The terms of the note require Stratton to pay 10 equal payments of $19,022 each December 31 for 10 years. The required general journal entry to record the payment on the nóte on December 31, Year 2 is: Debit Interest Expense $8400: debit Notes Payable $10,622; credit Cash $19,022 Debit Notes Payable $140,000; debit interest Expense $5022; credit Cash $19,022 Debit Interest Expense $7763 debit Notes Payable S 11,25% credit Cash S 19.022. Debit Interest Expense 58400; debit Notes Payable $10,622 credit Cash $19,022. Debit Notes Payable $8400; debit Interest Expense $10,622, credit Cash $19,022 Debit Notes Payable $19,022: credit Cash $19,022 Selected Answer: Answers:
Question 10 1 out of 1 points On anuary 1, a company issues bonds dated anuary 1 with a par value of S260,0 . The bonds mature in 3 years. The contract rate is 9.0%, and interest s paid semiannually on June 30 and December 31, The market rate is 10.0%. Using the present value factors below, the issue (selling) price of the bonds is: i 9.0% Present Value of an Annuity Present value of $ n- 2.5313 5.1579 2.4869 5.0757 0.7722 0.7679 0.7513 0.7462 64.9% 3, 10.0% 50% Selected Answer: s253,398 $194,012 $266,602 $260,000 $59,386 $253,398. Answers:
Question 11 1 out of 1 points On November 1 Alan Company signed a 120-day, 9% note payable with a face value of $66,000. Alan made the appropriate year-end accrual. What is the journal entry as of March 1 to record the payment of the note assuming no reversing entry was made? (Use 360 days a year.) Selected Answer: Debit Notes Payable $66,000, debit Interest Payable $990; debit Interest Expense $990; credit Cash $67,980. Answers: G Debit Notes Payable $66,000; debit Interest Payable $990; debit Interest Expense $990; credit Cash $67,980. Debit Notes Payable $67,980; credit Interest Payable $990; credit Interest Expense $990; credit Cash $66,000 Debit Cash $66,990; credit Notes Payable $66,990 Debit Notes Payable $66,000; debit Interest Expense $1980; credit Cash $67.980. Debit Notes Payable $66,000; debit Interest Payable $990; credit Cash s66,990. Question 12 t out of 1 points Peavey Enterprises purchased a depreciable asset for $24,000 on April 1, Year 1. The asset will be depreciated using the straight-Jine method over its four-year useful life. Assuming the assets salvage value is $2400, Peavey Enterprises should recognize depreciotion expense in Year 2 畛 in the amount of: Selected Answer:$5400.00 Answers: s540000 $6000.00 $20,700.00 $21,600.00 $4500.00
Question 13 1 out of 1 points Portia Grant is an employee who is paid monthly. For the month ofjanuary of the current year, she earned a total of 9188. The FICA tax for social security is 6.2% of the first $118,500 of employee earnings each calendar year and the FICA tax rate for Medicare is 1 45% of all earnings. The FUTA tax rate of 0.6% and the SUTA tax rate of SA% are applied to the first S7000 of an employees pay. The amount of federal income tax withheld from her earnings was $152457. Her net pay for the month is: (Round your intermediate calculations to two decimal places Selected Answer:$6960.54 $6526.54 $5041.43 $7530.20 $6566.00 О $6960,54 Answers: 1 out of 1 points Question 14
15 When originally purchased, a vehicle costing $23,940 had an estimated useful life of 8 years and an estimated salvage value of $2100. After 4 years of straight-line depreciation, the assets total estimated useful life was revised from 8 years to 6 years and there was no change in the estimated salvage value. The depreciation expense in year 5 equals $2898.00. $5628.00. $2730.00. $10,920.00. $5460.00. 23,940-200/2-22672:S
Question 5 0 out of 1 points A corporation issued 120 shares of its $5 par value common stock in payment of a $2000 charge frorn its accountant for assistance in filing its charter with the state. The entry to record this transaction will include: Selected Answer: A$2000 credit to Cash. Answers: A $2000 credit to Common Stock. A $600 debit to Organization Expenses. A $1400 credit to Paid-in Capital In Excess of Par Value, Common Stock A $2000 debit to Legal Expenses. A $2000 credit to Cash Question 6 0 out of 1 points g) If a check correctly written and paid by the bank for $246 is incorrectly recorded in the companys books for $264, how should this error be treated on the bank reconciliation? Selected Answer: ) Subtract %18 from the book balance. Subtract $18 from the banks balace and add $45 to the books balance Subtract $18 from the book balance. Add $18 to the book balance. Add $18 to the banks balance. Subtract $18 from the banks balance Answers
Answer #1

as per HomeworkLib policy i am to answer first 4 questions still i am answering first 5 questions , for rest kindly repost i am sure you will be helped

Ques 9
hence your answer debit interest epense 7763 ; notes payabe 11259 and credit cash 19022
Time Opening Balance EMI Intt @ 6% Principal (EMI-Intt) Closing Balnce
1 $                140,000 $   19,022 $     8,400 $                       10,622 $           129,378
2 $                129,378 $   19,022 $     7,763 $                       11,259 $           118,119
3 $                118,119 $   19,022 $     7,087 $                       11,935 $           106,184
4 $                106,184 $   19,022 $     6,371 $                       12,651 $             93,533
5 $                  93,533 $   19,022 $     5,612 $                       13,410 $             80,123
6 $                  80,123 $   19,022 $     4,807 $                       14,215 $             65,908
7 $                  65,908 $   19,022 $     3,954 $                       15,068 $             50,841
8 $                  50,841 $   19,022 $     3,050 $                       15,972 $             34,869
9 $                  34,869 $   19,022 $     2,092 $                       16,930 $             17,939
10 $                  17,939 $   19,022 $     1,076 $                       17,939 $                        0
Que 10
semi-annual interest = 260000 * 9%*6/12 = 11700
ao there will total 6 interest payments
present value of
always when there are semi annual payments 2 changes happen interest rate is halved and time is doubled
annual interest = 11700 * PVIFA (i=5% , t= 6years)=11700*5.0757= 59385.69
principal repayment = 260000 * PVIF (i=5% , t= 6 year)=260000*0.7513= 194012
total $           253,398
Ques 11
so on year end december 31 , you would have accrued
66000*9%*60/360 days = $         990 interest
so this would be recorded as laibilty in the books In the name of
"Interest payable"
so on march 1 since no reversing entry has been made , this account needs to be closed , because
the note will be paid off
so we close the interest payable account by debiting it
Ques 12
SLM = (Cost -Salvage value) / useful life
SLM=(24000-2400)/4= 5400
In year 2 also the same depriciation will be recognised since the straight line
method we assume the same depriciation every year
Ques 13
Gross Pay $              9,188.00
Social security $                  569.66 9188*6.2%
Medicare $                  133.23 9188*1.45%
witholding taxes $              1,524.57
Net pay $              6,960.55